# CLQCD Collaboration

### v2

@@ II Numerical setup
where $H$ represents an arbitrary hadronic state. $Z_{V}(H)$ should be independent of $H$ in the continuum due to charge conservation, while it can suffer from a significant $m_{q}a$ error at finite lattice spacing. In Fig. 3, we show that $Z_{V}(PS)$ is almost a single function of $m_{\rm PS}^{2}a^{2}$ and insensitive to lattice spacing, after it is normalized by $Z_{V}(\pi)$ in the chiral limit. Based on the joint fit, it can be approximated by the following relation,
$Z_{V}({\rm PS};a)=Z_{V}(0)\big{\{}1+(0.587(9)-0.43(6)a^{2})m_{\rm PS}^{2}a^{2}$
$\quad\quad-(0.131(8)-0.35(4)a^{2})m_{\rm PS}^{4}a^{4}+{\cal O}(m_{\rm PS}^{6}a^{6})\big{\}}.$ (9)
at 0.3% level.
Figure 4: $Z^{\rm RI^{\prime}}_{V}$ from the Landau gauge fixed propagators for the quark bilinear operators $\bar{l}l$ (red circles), $\bar{l}c$ (blue crosses), and $\bar{c}c$ (green boxes), at $a=0.1053$ fm. After the $a^{2}p^{2}$ extrapolation with $a^{2}p^{2}\geq 4$ shown as the curves, the extrapolated values agree with the chiral extrapolated value $Z_{V}(\pi)=0.79814(23)$ within 2%.
At the same time, one can also calculate $Z_{V}$ through the quark propagators directly,
$Z^{\rm RI^{\prime}}_{V}(a,a^{2}p^{2})=\frac{C_{V}Z_{q}}{\frac{1}{48}\text{Tr}[\Lambda^{\mu}_{V}(p)\gamma_{\mu}]},$ (10)
[…]
$C_{V}=\frac{\text{Tr}[\Lambda^{\mu}_{V}(p)\gamma_{\mu}]}{\text{Tr}[\Lambda^{\mu}_{V}(p)\gamma_{\nu}(g_{\mu\nu}-\frac{p_{\mu}p_{\nu}}{p^{2}})]}$ (11)
can be calculated perturbatively to avoid the discretization error in the off-diagonal projection $\text{Tr}[\Lambda^{\mu}_{V}(p)\gamma_{\nu\neq\mu}]$. As shown in Fig. 4, it is surprising that $Z^{\rm RI^{\prime}}_{V}$ for the quark bilinear operators $\bar{l}l$ (red circles), $\bar{l}c$ (blue crosses), and $\bar{c}c$ (green boxes) for the light quark $l$ and charm quark $c$ are similar within $\sim 1$% in the perturbative region $p^{2}\geq(5\mathrm{GeV})^{2}$, and their $a^{2}p^{2}$ extrapolated values based on the values in the perturbative region are also consistent with $Z_{V}$ in the chiral limit with 2%, as illustrated in Fig. 3 at the largest lattice spacing where $Z_{V}(\eta_{c})$ for the charm quark is around 2 times of $Z_{V}(\pi)$. Thus, the $a^{2}m_{\rm PS}^{2}$ dependence of $Z_{V}$ would be an additional discretization effect of the heavy quark in the hadron.
Fig. 5 displays the values of $f_{J/\psi}$ at various lattice spacings using either $Z_{V}^{c}\equiv Z_{V}(\eta_{c})$ (blue triangles) or $Z_{V}\equiv Z_{V}(\pi)_{m_{\pi}\rightarrow 0}$ (red circles). It is evident that employing the heavy quark improved normalization $Z_{V}^{c}$ leads to significantly smaller discretization errors and also better agreement with the experimental value 406.5(3.7)(0.5) MeV Workman et al. (2022); Hatton et al. (2020a) after the continuum extrapolation, compared to the case with the chiral extrapolated normalization $Z_{V}$.
The improvement in suppressing discretization errors for heavy flavor matrix elements can be extended to other currents as well. For instance, considering $g_{S,{\rm ME}}(H)\equiv\frac{\langle H|S|H\rangle}{\langle H|H\rangle}$ as illustrated in Fig. 6, one can use the original $Z_{S}$ (red circles) for the $\bar{c}c$ matrix element in the $\eta_{c}$, akin to the treatment for $\bar{l}l$ in the pion, or opt for the heavy quark improved $Z_{S}^{c}\equiv Z_{V}^{c}\frac{Z_{S}}{Z_{V}}$ (blue triangles) to mitigate discretization errors. Additionally, we present $g_{S,{\rm FH}}\equiv\frac{\partial m_{\eta_{c}}}{\partial m^{R}_{c}}=\frac{Z_{S}}{Z_{A}}\frac{\partial m_{\eta_{c}}}{\partial m^{\rm PCAC}_{c}}$ (green crosses) for comparison, which should align with $g_{S,{\rm ME}}$ in the continuum. Upon continuum extrapolation, all these quantities converge, while $g_{S,{\rm ME}}$ renormalized by $Z_{S}$ exhibits a larger discretization error.
The improvement can be extended to the flavor-changed current, like $X\equiv\bar{c}\Gamma l$, by utilizing the improved renormalization constant $Z_{X}^{c}\equiv\sqrt{Z_{V}Z_{V}^{c}}\frac{Z_{X}}{Z_{V}}$. As shown in Fig. 7, for the decay constant of $D_{s}^{*}$, we can see that the use of the heavy quark improved renormalization constant can reduce the discretization error significantly, while the continuum limit from both cases agree with each other. Similar improvement could be used for decay constant of the other open-charm meson, such as $f_{D}$, as discussed in the following section.
As shown in Eq. (II), the $O(a^{6}m_{\rm PS}^{2})$ term is necessary to describe the $m_{\rm PS}^{2}a^{2}$ dependence of $Z_{V}({\rm PS};a)$. Thus the differences between the $f_{J/\psi}$ and $g_{S,\eta_{c}}$ results using $Z_{V}$ and $Z_{V}^{c}$ are the ${\cal O}(a^{6})$ effect which is not included in the current fit ansatz, especially the cases using the chiral extrapolated $Z_{V}$ in the charmoninum matrix elements. Since using $Z_{V}^{c}$ can significantly suppress the discretization error, we will only use $Z_{V}^{c}$ for all the following discussion when the charm quark is involved.
Figure 7: Comparison of $f_{D_{s}^{*}}$ using heavy quark improved normalization constant $Z_{V}^{cl}\equiv\sqrt{Z_{V}Z_{V}^{c}}$ and original normalization constants $Z_{V}$ at different lattice spacing and continuum. The continuum limits obtained for both cases are consistent with each other, but the discretization error for the former is smaller.
$Z_{V}(a,m_{\pi},m_{\eta_{s}},1/L)=Z_{V}(a,0,0,0)$
and re-scale the uncertainty by $\sqrt{\chi^{2}/\mathrm{d.o.f}}$. All the chiral-extrapolated renormaralzation constants are collected in Table 2 and the numbers on each ensemble without chiral extrapolation can be found in Table 4 in the appendix. It can be observed that $Z_{V}^{c}$ is a factor of 2 larger than $Z_{V}$ at the coarsest lattice spacing, while becomes much closer at the finest one.
$a$(fm) $Z_{V}$ $Z_{V}^{c}$ $Z_{A}/Z_{V}$ $Z_{S}/Z_{V}$ $Z_{P}/Z_{V}$ $Z_{T}/Z_{V}$
0.10530(18) 0.8001(05) 1.5744(17) 1.0728(04) 1.2013(31) 0.9239(48) 1.0816(06)
0.08877(30) 0.8203(07) 1.4042(20) 1.0561(12) 1.0948(75) 0.8640(99) 1.0958(18)
0.07750(18) 0.8365(05) 1.3071(13) 1.0537(05) 1.0487(30) 0.8428(39) 1.1021(07)
0.06826(27) 0.8478(06) 1.2392(14) 1.0423(06) 1.0052(74) 0.8149(47) 1.1101(13)
0.05187(26) 0.8703(06) 1.1321(15) 1.0353(08) 0.9162(95) 0.7842(62) 1.1238(14)
@@ III Results
Similar to our previous work Hu et al. (2024), we perform bootstrap re-samplings on each ensemble and conduct the correlated global fit based on these bootstrap samples. The lattice spacing and renormalization constants are also sampled for each bootstrap sample using a Gaussian distribution with their uncertainties as the width of the distribution. Thus our results presented in this section have included the systematic uncertainties from the lattice spacing determination and non-perturbative renormalization.
@@
where $m_{\pi}^{\rm phys}=134.98$ MeV is the physical $\pi^{0}$ mass, and the squares of the pion and $\eta_{s}$ masses are used to account for the dependence on the light quark mass and strange sea quark masses (as the strange valence quark mass has been interpolated to its physical value), respectively. The $a^{2}+a^{4}$ corrections are also introduced to describe the discretization error, and we will see that the $a^{4}$ term is essential for certain cases, especially the charm quark mass under $\overline{\textrm{MS}}$ 2 GeV. The $a\alpha_{s}$ term is not taken into account as it should be mainly removed by the non-perturbative renormalization, and our previous study Hu et al. (2024) suggests that $a^{2}$ extrapolation works well for light flavor physics. Based on the joint fit, one can further eliminate the impact from the unphysical light and strange quark masses by subtracting the $d^{X}_{1,2}$ terms from $X(m_{\pi}),m_{\eta_{s}},a)$ and obtain $X(m_{\pi}^{\rm phys},m_{\eta_{s}}^{\rm phys},a)$ with physical light and strange quark masses but at a finite lattice spacing.
@@
$m^{\rm QCD}_{D}$ $=1869.2(4.0)~{}\mathrm{MeV},$
$m^{\rm QCD}_{D^{\pm}}-m^{\rm QCD}_{D^{0}}$ $=d^{D}_{1}\frac{(m_{\pi}^{\rm phys})^{2}}{m_{l}}(m_{d}-m_{u})+{\cal O}(\frac{m_{l}^{2}}{\Lambda_{\chi}})$
$=2.87(26)~{}\mathrm{MeV},$ (16)
@@
At the same time, our prediction of the $D^{*}$ mass (light blue) and the $D_{s}^{*}$ mass (green) are consistent with the experimental value. The unknown QED effect of the vector meson masses can be significant if the statistics uncertainty can be further suppressed.
@@
$\Delta^{\rm QCD,conn}_{{\rm HFS},\bar{c}c}=m_{J/\psi}^{\rm QCD,conn}-m_{\eta_{c}}^{\rm QCD,conn}=118.0(2.8)~{}\mathrm{MeV},$ (17)
and consistent with the previous HPQCD study Hatton et al. (2020b) using staggered fermions, 118.6(1.1) MeV. Such a value is slightly larger than the experimental value 113.0(5) Workman et al. (2022). The difference can be attributed mostly to the quark annihilation effect (The correction from the QED effect is around 1 MeV Hatton et al. (2020b)). Direct calculations using anisotropic $N_{f}$=2 charm quark sea ensemble at $a_{t}=0.0205$ fm with $a_{s}/a_{t}=5$ indicate that the inclusion of charm quark annihilation diagram almost does not change the $J/\psi$ mass, but pushes the $\eta_{c}$ mass up by approximately 3-4 MeV Zhang et al. (2022a, b), regardless of whether the mixing with the pseudo-scalar glueball is considered.
@@
which has been verified by the experiment at 0.1 MeV level and can be explained by the short range feature of the spin-spin interaction in the context of potential models Kwong et al. (1987); Lucha et al. (1991), we skipped the calculation of the $\chi_{c2}$ mass which requires additional complication in the interpolating field, and extract the following three splittings based on $m_{\chi_{c0},\chi_{c1},h_{c}}$ we obtained,
@@
$=15.7(3.8)~{}\mathrm{MeV}.$ (19)
All of three splittings agree with the experimental value and previous LQCD calculation using the HISQ fermion Burch et al. (2010) well within the uncertainty.
From Fig. 8, we can also see that the pseudo-scalar meson masses have much smaller lattice spacing dependence than the vector meson masses, since our valence strange and charm quark masses are also determined using the pseudo scalar masses. The lattice spacing dependence of the vector meson masses also suggests that the hyperfine splitting $\Delta_{\rm HFS}$ will exhibit a significant variation, which is crucial for ensuring that the continuum extrapolated values of $\Delta_{\rm HFS}$ closely match the physical values.
@@
Using the charm quark improved normalization, the decay constants of the pseudo-scalar and vector open-charm mesons and charmonium at different lattice spacings are illustrated in Fig. 9. The decay constants of the open-charm mesons exhibit negligible lattice spacing dependence and are consistent with the previous HPQCD results using the HISQ action within a 1-2% uncertainty. For the values of $f_{J\psi}$ and $f_{\eta_{c}}$, after using the charm quark improved normalization, the lattice spacing dependence is weakened and our final predictions are consistent with that from HPQCD.
Figure 10: Similar to Fig. 8 but for the transverse decay constants.
Figure 11: Similar to Fig. 8 but for the P-wave charmonium decay constants.
In addition, we show the lattice spacing dependence of transverse decay constant of the charmed vector mesons in Fig. 10 and also those of the P-wave charmonium in Fig. 11.
@@
Using a similar approach as outlined in Ref. Hu et al. (2024) for the light and strange quarks, the renormalized charm quark mass at $\overline{\textrm{MS}}$ 2 GeV with $N_{f}=3$ matching and running is determined for each ensemble. One can further obtain the charm quark mass in the continuum and also physical light and strange quark mass by using the joint fit the data on different ensembles. After correcting for small effects arising from the unphysical light and strange quark masses, the renormalized charm quark masses on all the ensembles with different lattice spacings are depicted in Fig. 12. It is evident that at the largest lattice spacing, $m_{c}$ deviates by $\sim$10from its continuum extrapolated value, based on a $a^{2}+a^{4}$ extrapolation. Converting the charm quark mass to $m_{c}(m_{c})$ using perturbative running yields a value of 1.289(17) GeV, with systematic uncertainties from lattice spacing determination and renormalization.
Table 3: Charm quark mass under the $\overline{\textrm{MS}}$ scheme, charmed meson masses, and also related decay constants. The so-call “Pure QCD” values corresponds to the values from literatures without annihilation diagram or QED correction. All the values are in unit of GeV.
@@
This work 1.289(17)(01) 1.8692(40)(01) 2.0162(49)(10) 2.1130(54)(02)
FLAG/Exp. 1.276(05) 1.8672(2) 2.00856(5) 2.1122(04)
@@
This work 0.2102(33)(03) 0.2321(43)(01) 0.2073(39)(08) 0.2487(28)(02) 0.2743(34)(01) 0.2449(47)(01)
Literature 0.2090(24) Aoki et al. (2022) 0.234(6) Chen et al. (2021) 0.192(14) Li et al. (2024) 0.2480(16) Aoki et al. (2022) 0.274(6) Donald et al. (2014) 0.252(10) Li et al. (2024)
@@
This work 0.3981(44)(14) 0.4059(51)(25) 0.3923(53)(18) 0.314(23)(01) 0.2105(91)(03) 0.1464(80)(03)
Literature 0.3975(10) Hatton et al. (2020a) 0.4096(16) Hatton et al. (2020a) 0.3911(56) Li et al. (2024)
Eventually we just collect our predictions in Table 3 for comparison. Using the results of $f_{D}$ and $f_{D_{s}}$, one can extract the CKM matrix elements $|V_{cd}|$ and $|V_{cs}|$, respectively. Taking into account the experimental constraints on the products of decay constants and CKM matrix elements Workman et al. (2022), as given below
$f_{D^{+}}|V_{cd}|$ $=$ $45.8(1.1)\textrm{MeV}$
$f_{D_{s}^{+}}|V_{cs}|$ $=$ $243.5(2.7)\textrm{MeV}$ (20)
@@
$|V_{cd}|=0.2179(33)(52),\quad|V_{cs}|=0.979(11)(11),$ (21)
where the first errors come from the lattice results and the second result from the uncertainties from the experiments in Eq. (III). The above results are consistent with those quoted by the PDG, but with slight larger errors. Using the present PDG value of $|V_{cb}|=0.041(1)$ Workman et al. (2022) which is negligible comparing to the uncertainties of $|V_{cd(s)}|$, we predict,
$|V_{cd}|^{2}+|V_{cs}|^{2}+|V_{cb}|^{2}=1.008(23)(23),$ (22)
@@
Our predictions of the quantities $f_{D^{*}}$ and $f_{D_{s}^{*}}$ are by far the most precise. Comparing with the previous resutls Chen et al. (2021); Donald et al. (2014); Li et al. (2024), our calculation provides much better control on the systematic uncertainties on the unphysical sea quark masses and discretization error, and then also by far the most accurate. Therefore, we can provide insight into the experimental search for the purely leptonic decay of $D^{*}$ and $D_{s}^{*}$ particles. Most recently, the BESIII collaboration has reported the first experimental measurements on the purely leptonic decays $D^{*}\rightarrow l\nu_{l}$ Ablikim et al. (2024) and $D_{s}^{*}\rightarrow l\nu_{l}$ Ablikim et al. (2023) with $l=e,\mu$. No significant signal is observed for the channel $D^{*}\rightarrow l\nu_{l}$, and only upper limits are set as $\operatorname{Br}(D^{*}\rightarrow e\nu_{e})<1.1\times 10^{-5}$ and $\operatorname{Br}(D^{*}\rightarrow\mu\nu_{\mu})<4.3\times 10^{-6}$. For $D_{s}^{*}\rightarrow l\nu_{l}$, the BESIII gives the branching fraction of this decay as $(2.1^{+1.2}_{-0.9_{\textrm{stat.}}}\pm 0.2_{\textrm{syst.}})\times 10^{-5}$. Theoretically, the decay widths of these pure leptonic decays can be parameterized by the $f_{D^{*}}$ and $f_{D_{s}^{*}}$ via the following way
$\Gamma=\frac{G_{F}^{2}}{12\pi}|V_{cd(s)}|f_{V}^{2}m_{V}^{3}\left(1-\frac{m_{l}^{2}}{m_{V}^{2}}\right)^{2}\left(1+\frac{m_{l}^{2}}{2m_{V}^{2}}\right)$ (23)
@@
$\Gamma(D^{*}\rightarrow l\nu_{l})_{l=e,\mu}$ $=$ $3.54(16)\times 10^{-7}~{}\mathrm{keV},$
$\Gamma(D_{s}^{*}\rightarrow l\nu_{l})_{l=e,\mu}$ $=$ $2.41(09)\times 10^{-6}~{}\mathrm{keV},$ (24)
The branching fractions of them can then be determined by combining with the total decay width $\Gamma_{\textrm{total}}(D^{*})=83.4(1.8)~{}\textrm{keV}$ Workman et al. (2022) and $\Gamma_{\textrm{total}}(D_{s}^{*})=0.0587(54)~{}\textrm{keV}$ Meng et al. (2024). The results are given as $\operatorname{Br}(D^{*}\rightarrow l\nu_{l})=4.24(09)(19)\times 10^{-9}$ and $\operatorname{Br}(D_{s}^{*}\rightarrow l\nu_{l})=4.11(38)(15)\times 10^{-5}$.

### v3

@@ II Theoretical framework
where $H$ represents an arbitrary hadronic state. $Z_{V}(H)$ should be independent of $H$ in the continuum due to charge conservation, while it can suffer from a significant quark mass dependence at finite lattice spacing. As shown in Ref. Capitani et al. (2001) and verified by our practical calculation, this dependence majorly comes from the massive quark propagator on the lattice and can be absorbed into the redefinition of the massive quark field.
Thus we propose to use the heavy quark improved normalization factor $Z_{V}^{c}\equiv Z_{V}(\eta_{c})$ for the $\bar{c}\gamma_{\mu}c$ current, and define that of the other charm quark bilinear current $X$ through the chiral-extrapolated ratio $\frac{Z_{X}}{Z_{V}}$ under the $\overline{\mathrm{MS}}$ scheme through the intermediate RI/MOM scheme Martinelli et al. (1995), $Z_{X}^{c}\equiv Z_{V}^{c}\frac{Z_{X}}{Z_{V}}$. The improvement can be extended to the flavor-changed current, like $Y\equiv\bar{c}\Gamma l$, by utilizing the improved renormalization constant $Z_{Y}^{c}\equiv\sqrt{Z_{V}Z_{V}^{c}}\frac{Z_{Y}}{Z_{V}}$. Thus the improvement here can be understood by a redefinition of the charm quark field normalization at finite $a$. Based on the numerical tests on $f_{J/\psi}$, $f_{D^{*}_{s}}$ and $g_{S,{\rm ME}}(H)\equiv\frac{\langle H|S|H\rangle}{\langle H|H\rangle}$, this heavy quark improved renormalization can suppress the discretization error of heavy quark significantly, as detailed in Sec. C of the supplemental materials. The values of the renormalization constants, and also the lattice spacing dependence of the transverse decay constant of the charmed vector mesons and those of the P-wave charmonium can also be found there.
@@ III Results
Similar to our previous work Hu et al. (2024), we perform bootstrap re-samplings on each ensemble and conduct the correlated global fit based on these bootstrap samples. Since we require the lattice spacing $a$ and $Z_{V}$ to be the same with similar $\beta$ and to correspond to physical light and strange quark masses, $a$ and $Z_{V}$ are obtained as the fit parameters through joint fits using kinds of quark masses and $\beta$. Then the bootstrap samples are generated for $a$ and $Z_{V}$ to keep the correlation between different ensembles at the same $\beta$. This treatment ensures the statistical uncertainty of $a$ and $Z_{V}$ contribute to that of our final predictions properly.
@@
where $m_{\pi}^{\rm phys}=134.98$ MeV is the physical $\pi^{0}$ mass, and the squares of the pion and $\eta_{s}$ masses are used to account for the dependence on the light quark mass and strange sea quark masses (as the strange valence quark mass has been tuned to its physical value), respectively. The $a^{2}+a^{4}$ corrections are also introduced to describe the discretization error, and we will see that the $a^{4}$ term is essential for certain cases, especially the charm quark mass under $\overline{\textrm{MS}}$ 2 GeV. The $a\alpha_{s}$ term is not taken into account as it should be mainly removed by the non-perturbative renormalization, and our previous study Hu et al. (2024) suggests that $a^{2}$ extrapolation works well for light flavor physics. Based on the joint fit, one can further eliminate the impact from the unphysical light and strange quark masses by subtracting the $d^{X}_{1,2}$ terms from $X(m_{\pi},m_{\eta_{s}},a)$ and obtain $X(m_{\pi}^{\rm phys},m_{\eta_{s}}^{\rm phys},a)$ with physical light and strange quark masses but at a finite lattice spacing.
@@
The systematic uncertainty from the finite volume correction is estimated by including additional $e^{-m_{\pi}L}$ term in the joint fit and the impact turns out to be negligible compared to the statistical uncertainty. A similar treatment is used to estimate the systematic uncertainty from the possible ${\cal O}(a\alpha_{s})$ term, which is also small.
Since we used FLAG average value of the scale parameter $w_{0}$=0.1736(9) fm Aoki et al. (2022) when determining the lattice spacing, this error of $w_{0}$ will enter the final result as system uncertainty. To estimate this effect, we vary the central value of $w_{0}$ by 1 sigma, redo the entire analysis, and consider the difference in the central value of the result as a systematic uncertainty. Since we determine the bare charm quark mass parameter on each ensemble using the physical $D_{s}$ mass, the impact of the $w_{0}$ uncertainty on the charmed hadron mass will be partially canceled and then can be smaller than 0.5%.
@@
$m^{\rm QCD}_{D}$ $=1863.7(1.8)(1.5)~{}\mathrm{MeV},$
$m^{\rm QCD}_{D^{\pm}}-m^{\rm QCD}_{D^{0}}$ $=d^{D}_{1}\frac{(m_{\pi}^{\rm phys})^{2}}{m_{l}}(m_{d}-m_{u})(1+{\cal O}(\frac{(m_{\pi}^{\rm phys})^{2}}{\Lambda^{2}_{\chi}}))$
$=2.88(22)~{}\mathrm{MeV},$ (8)
@@
At the same time, our prediction of the $D^{*}$ mass (light blue) and the $D_{s}^{*}$ mass (green) are consistent with the experimental value. The unknown QED effect of the vector meson masses can be significant if the statistical uncertainty can be further suppressed.
@@
$\Delta^{\rm QCD,conn}_{{\rm HFS},\bar{c}c}$ $=m_{J/\psi}^{\rm QCD,conn}-m_{\eta_{c}}^{\rm QCD,conn}$
$=115.3(1.4)(1.1)~{}\mathrm{MeV},$ (9)
and slightly smaller than the previous HPQCD study Hatton et al. (2020a) using staggered fermions, 118.6(1.1) MeV. Such a value is slightly larger than the experimental value 112.8(4) Workman et al. (2022). The difference can be attributed mostly to the quark annihilation effect (The correction from the QED effect is around 1 MeV Hatton et al. (2020a)). Direct calculations using anisotropic $N_{f}$=2 charm quark sea ensemble at $a_{t}=0.0205$ fm with $a_{s}/a_{t}=5$ indicate that the inclusion of charm quark annihilation diagram almost does not change the $J/\psi$ mass, but pushes the $\eta_{c}$ mass up by approximately 3-4 MeV Zhang et al. (2022a, b), regardless of whether the mixing with the pseudo-scalar glueball is considered.
@@
which has been verified by the experiment at 0.1 MeV level and can be explained by the short-range feature of the spin-spin interaction in the context of potential models Kwong et al. (1987); Lucha et al. (1991), we skipped the calculation of the $\chi_{c2}$ mass which requires additional complication in the interpolating field, and extract the following three splittings based on $m_{\chi_{c0},\chi_{c1},h_{c}}$ we obtained,
@@
$=15.4(7.5)(0.2)~{}\mathrm{MeV}.$ (11)
All three splittings agree with the experimental value and previous LQCD calculation using the HISQ fermion Burch et al. (2010) well within the uncertainty.
From Fig. 2, we can also see that the pseudo-scalar meson masses have much smaller lattice spacing dependence than the vector meson masses since our valence strange and charm quark masses are also determined using the pseudoscalar masses. The lattice spacing dependence of the vector meson masses also suggests that the hyperfine splitting $\Delta_{\rm HFS}$ will exhibit a significant variation, which is crucial for ensuring that the continuum extrapolated values of $\Delta_{\rm HFS}$ closely match the physical values.
@@
Using the charm quark improved normalization, the decay constants of the pseudo-scalar and vector open-charm mesons and charmonium at different lattice spacings are illustrated in Fig. 3. The decay constants of the open-charm mesons exhibit negligible lattice spacing dependence and are consistent with the previous HPQCD results using the HISQ action within a 1-2% uncertainty. For the values of $f_{J/\psi}$ and $f_{\eta_{c}}$, after using the charm quark improved normalization, the lattice spacing dependence is weakened and our final predictions are consistent with that from HPQCD.
@@
Using a similar approach as outlined in Ref. Hu et al. (2024) for the light and strange quarks, the renormalized charm quark mass at $\overline{\textrm{MS}}$ 2 GeV with $N_{f}=3$ matching and running is determined for each ensemble. One can further obtain the charm quark mass in the continuum and also physical light and strange quark mass by using the joint fit of the data on different ensembles. After correcting for small effects arising from the unphysical light and strange quark masses, the renormalized charm quark masses on all the ensembles with different lattice spacings are depicted in Fig. 4. It is evident that at the largest lattice spacing, $m_{c}$ deviates by $\sim$10from its continuum extrapolated value, based on a $a^{2}+a^{4}$ extrapolation. Converting the charm quark mass to $m_{c}(m_{c})$ using perturbative running results in a value of 1.2933(72)(95) GeV, where the two uncertainties represent the statistical and systematic errors, respectively.
Table 2: Charm quark mass under the $\overline{\textrm{MS}}$ scheme, charmed meson masses, and also related decay constants. Note that our predictions and those from the literature do not include the contribution from either the annihilation diagram or QED correction. All the values are in the unit of GeV, and all the Experimental values come from PDG Workman et al. (2022).
@@
This work 1.2933(72)(95) 1.8637(18)(15) 2.0115(25)(28) 2.1130(22)(22)
FLAG/Exp. 1.276(05) Aoki et al. (2022); McNeile et al. (2010); Yang et al. (2015); Nakayama et al. (2016); Petreczky and Weber (2019); Heitger et al. (2021) 1.8672(02) 2.0086(01) 2.1122(04)
@@
This work 0.2108(20)(11) 0.2292(26)(17) 0.2068(27)(14) 0.2470(21)(04) 0.2691(30)(03) 0.2447(34)(06)
Literature 0.2084(15) Kuberski et al. (2024) 0.234(6) Chen et al. (2021) 0.192(14) Li et al. (2024) 0.2468(13) Kuberski et al. (2024) 0.274(6) Donald et al. (2014) 0.252(10) Li et al. (2024)
@@
This work 0.3944(29)(18) 0.4114(34)(22) 0.3893(32)(24) 0.314(17)(04) 0.2103(79)(34) 0.1465(73)(31)
Literature 0.3975(10) Hatton et al. (2020b) 0.4096(16) Hatton et al. (2020b) 0.3911(56) Li et al. (2024)
Eventually, we just collect our predictions in Table.  2 for comparison. Besides systematic uncertainties mentioned above, those from the experimental value of $m_{D_{s}}$ and perturbative matching are not negligible. For the details of the systematic uncertainties estimation and error budget, see Sec. D of Supplementary Material.
Using the results of $f_{D}$ and $f_{D_{s}}$, one can extract the CKM matrix elements $|V_{cd}|$ and $|V_{cs}|$, respectively. Taking into account the experimental constraints on the products of decay constants and CKM matrix elements Workman et al. (2022), as given below
$f_{D^{+}}|V_{cd}|$ $=$ $\ \ \!45.8(1.1)~{}\textrm{MeV},$
$f_{D_{s}^{+}}|V_{cs}|$ $=$ $243.5(2.7)~{}\textrm{MeV},$ (12)
@@
$|V_{cd}|=0.2173(21)_{\rm lat}(52)_{\rm ex},\quad|V_{cs}|=0.986(08)_{\rm lat}(11)_{\rm ex},$ (13)
where the first errors come from the lattice results and the second result from the uncertainties from the experiments in Eq. (III). The above results are consistent with those quoted by the PDG, but with slightly larger errors. Using the present PDG value of $|V_{cb}|=0.041(1)$ Workman et al. (2022) which is negligible comparing to the uncertainties of $|V_{cd(s)}|$, we predict,
$|V_{cd}|^{2}+|V_{cs}|^{2}+|V_{cb}|^{2}=1.021(17)(22),$ (14)
@@
Based on previous CLQCD work Hu et al. (2024) on similar ensembles, one can extract $f_{K}/f_{\pi}=1.1907(76)_{\rm stat}(17)_{\rm sys}$, where the systematic uncertainty from the renormalization is largely canceled out. By integrating this result with the experimental constraint $\frac{|V_{us}|}{|V_{ud}|}\frac{f_{K}}{f_{\pi}}=0.27683(29)_{\rm exp}(20)_{\rm th}$ Workman et al. (2022). and the unitarity constraint is expressed as
$1=|V_{ud}|^{2}+|V_{us}^{2}|+|V_{ub}|^{2}=|V_{ud}|^{2}+|V_{us}^{2}|+0.0035^{2},$ (15)
we derive the values
$|V_{ud}|=0.9740(03)_{\rm lat}(01)_{\rm ph},\ |V_{us}|=0.2265(14)_{\rm lat}(03)_{\rm ph}.$ (16)
Consequently, the CKM relations $|V_{ud}|=|V_{cs}|$ and $|V_{us}|=|V_{cd}|$ are substantiated.
Our predictions of the quantities $f^{(T)}_{D^{*}}$ and $f^{(T)}_{D_{s}^{*}}$ are by far the most precise. Compared with the previous resutls Chen et al. (2021); Donald et al. (2014); Li et al. (2024), our calculation provides much better control on the systematic uncertainties on the unphysical sea quark masses and discretization error, and then also by far the most accurate. Therefore, we can provide insight into the experimental search for the purely leptonic decay of $D^{*}$ and $D_{s}^{*}$ particles. Most recently, the BESIII collaboration has reported the first experimental measurements on the purely leptonic decays $D^{*}\rightarrow l\nu_{l}$ Ablikim et al. (2024) and $D_{s}^{*}\rightarrow l\nu_{l}$ Ablikim et al. (2023) with $l=e,\mu$. No significant signal is observed for the channel $D^{*}\rightarrow l\nu_{l}$, and only upper limits are set as $\operatorname{Br}(D^{*}\rightarrow e\nu_{e})<1.1\times 10^{-5}$ and $\operatorname{Br}(D^{*}\rightarrow\mu\nu_{\mu})<4.3\times 10^{-6}$. For $D_{s}^{*}\rightarrow l\nu_{l}$, the BESIII gives the branching fraction of this decay as $(2.1^{+1.2}_{-0.9_{\textrm{stat.}}}\pm 0.2_{\textrm{syst.}})\times 10^{-5}$. Theoretically, the decay widths of these pure leptonic decays can be parameterized by the $f_{D^{*}}$ and $f_{D_{s}^{*}}$ via the following way
$\Gamma=\frac{G_{F}^{2}}{12\pi}|V_{cd(s)}|f_{V}^{2}m_{V}^{3}\left(1-\frac{m_{l}^{2}}{m_{V}^{2}}\right)^{2}\left(1+\frac{m_{l}^{2}}{2m_{V}^{2}}\right),$ (17)
@@
$\Gamma(D^{*}\rightarrow l\nu_{l})_{l=e,\mu}$ $=$ $3.388(98)\times 10^{-7}~{}\mathrm{keV},$
$\Gamma(D_{s}^{*}\rightarrow l\nu_{l})_{l=e,\mu}$ $=$ $2.432(83)\times 10^{-6}~{}\mathrm{keV},$ (18)
The branching fractions of them can then be determined by combining with the total decay width $\Gamma_{\textrm{total}}(D^{*})=83.4(1.8)~{}\textrm{keV}$ Workman et al. (2022) and $\Gamma_{\textrm{total}}(D_{s}^{*})=0.0587(54)~{}\textrm{keV}$ Meng et al. (2024a). The results are given as $\operatorname{Br}(D^{*}\rightarrow l\nu_{l})=4.06(09)(12)\times 10^{-9}$ and $\operatorname{Br}(D_{s}^{*}\rightarrow l\nu_{l})=4.08(38)(10)\times 10^{-5}$.
Our major results are summarized in Table 2, where the first error encompasses the statistical uncertainties from the correlation function, lattice spacing determination, pion and $\eta_{s}$ masses in the sea, and also non-perturbative renormalization. The second error includes the systematic uncertainty stemming from the fit parameterization, the FLAG scale parameter $w_{0}$, finite volume correction, missing ${\cal O}(a\alpha_{s})$ effects, and also perturbative matching of the renormalization. The values from PDG Workman et al. (2022), FLAG Aoki et al. (2022), HPQCD Hatton et al. (2020b); Donald et al. (2014) and ALPHA collaborations Kuberski et al. (2024) are also listed there for comparison. We use the values of $f_{D_{(s)}}$ Bussone et al. (2024); Kuberski et al. (2024) from the ALPHA collaboration instead of the FLAG averages Aoki et al. (2022); Davies et al. (2010); Bazavov et al. (2012); Na et al. (2012); Yang et al. (2015); Boyle et al. (2017) since they are slightly more precise. RBC/UKQCD also obtained high-precision values of $f_{D_{s}}/f_{D}=1.1740(51)(68)$ Boyle et al. (2018) and $m_{c}(m_{c})=1.292(12)$ GeV (using the heavy-quark improved renormalization) Del Debbio et al. (2024) which are consistent with our determination $f_{D_{s}}/f_{D}=1.1712(85)(41)$ and $m_{c}(m_{c})=1.2933(72)(95)$ GeV, but not included in the FLAG averages yet.